Use the given information to find the number of degrees of​ freedom, the critical values chi Subscript Upper L Superscript 2 and chi Subscript Upper R Superscript 2​, and the confidence interval estimate of sigma. It is reasonable to assume that a simple random sample has been selected from a population with a normal distribution. Platelet Counts of Women 90​% ​confidence; nequals21​, sequals65.4.
Answer #1

Solution :

Given that,

s = 65.4

s2 = 4277.16

n = 21

df = 20

\chi2L = \chi 2\alpha/2,df = 31.410

\chi2R = \chi 21 - \alpha /2,df = 10.851

The 90% confidence interval for \sigma is,

\sqrt(n - 1)s2 / \chi 2\alpha/2 < \sigma < \sqrt (n - 1)s2 / \chi 21 - \alpha /2

  \sqrt20 * 4277.16 / 31.410 < \sigma < \sqrt 20 * 4277.16 / 10.851

52.1862 < \sigma < 88.7896

(52.1862 , 88.7896 )